2.1 Consider a closed economy that is described by the following model:



C = 280m + 0.72Y



Where:



C = Consumption Y = Income



I = 150m I = Investment



G = 300m G = Government spending



T = 22% t = Tax rate.




2.1.1 Calculate the multiplier. (3)



2.1.2 Calculate the total autonomous spending. (2)


2.1.3 Calculate equilibrium Income(2)


2.1.4 Calculate the Level of savings at equilibrium(3)


2.1.5 Calculate the amount of tax collected at equilibrium(3)


1
Expert's answer
2022-03-21T12:53:00-0400

2.1.1 Calculate the multiplier. (3)

"\alpha" = 1/(1-C)

C(CMPC)="\Delta" C/"\Delta" Y =0.27= C

"\alpha" = 1/(1-0.72)

=3.5714


2.1.2 Calculate the total autonomous spending. (2)

"\Alpha\omicron" = C + I + G

= 280m + 150m +300m

= 730m

2.1.3 Calculate the equilibrium income. (3)


Y = C + I + G

Y = 280 + 0.72Y +150 + 300

Y - 0.72Y = 730

0.28Y = 730

0.28Y/0.28 = 730/0.28

Y= 2607.14M

Therefore equilibrium income is 2607.14M


2.1.4 Calculate the level of savings at equilibrium. (2)

C = 280 + 0.72Y

= 280 + 0.72(2607.14)

= 2157.15M

Y = C + S

2607.14 =2157.14 + S

S = 2607.14 - 2157.14

=450M


2.1.5 Calculate the amount of tax collected at equilibrium. (3)

tax rate is 22 % = 0.22

t(Y) = 0.22(2607.14)

= 573.5708M